multiply, (5sqrt(7)-15sqrt(2)) (4sqrt(7)+12sqrt(2))
Answer by ilana(307)

Use the FOIL method to multiply this out. Remember, (a*rtb) * (c*rtd) = ac*rtbd, so 5*rt7 * 4*rt7 = 20*rt49, or 20*7.
So here is the simplification:
20*7 + 60*rt14 - 60*rt14 - 180*2
20*7 - 180*2
140 - 360
-220